# Diagonalization $of$ Polynomial-Time Deterministic Turing Machines via Nondeterministic Turing Machines

### v31

@@ 7.1 Proof of Theorem 6
Now, we are ready to prove the Theorem 6:
@@
Next, we construct a five-tape universal nondeterministic oracle Turing machine $M_{0}^{O}$ which operates as follows on an input string $x$ of length of $n$:
  1. 1.
    $M^{O}_{0}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some polynomial-time deterministic oracle Turing machine $D^{O}_{j}$ for some $j$ then GOTO $5$, else determines $t$, the number of tape symbols used by $D^{O}_{j}$; $s$, its number of states; and $k$, its order.282828 We suppose that the order of $D^{O}$, i.e., the minimal degree of some polynomial of $D^{O}$, is also encoded into $D^{O}$, similarly to the way presented in Section 3. The third tape of $D^{O}_{0}$ can be used as “scratch” memory to calculate $t$.
  2. 2.
@@
    cells each, the blocks being separated by single cell holding a marker ##\##, i.e., there are
@@
  4. 4.
    On tape $4$, $M^{O}_{0}$ reads and writes the contents of oracle tape of $D^{O}_{j}$. That is, tape $4$ is the oracle tape of $M^{O}_{0}$ which is used to simulate the oracle tape of $D^{O}_{j}$.
  5. 5.
@@
  6. 6.
    Since $x$ is not encoding of some polynomial-time deterministic oracle Turing machine with oracle $O$. Then $M^{O}_{0}$ sets up a block of
@@
The nondeterministic oracle Turing machine $M^{O}_{0}$ described above is a nondeterministic oracle Turing machine which is of time complexity
@@
Suppose now $L^{O}_{d}$ were accepted by $i$-th deterministic oracle Turing machine $D^{O}_{i}$ in the enumeration $e$ which is a deterministic $T(n)=n^{k}+k$ time-bounded oracle Turing machine. Let $D^{O}_{i}$ have $s$ states and $t$ tape symbols. Since 292929 See (III) of footnote $6$.
@@
So, there exists a $N_{0}>0$ such that for any $N\geq N_{0}$,
$T(N)\log T(N)<N^{k+1}$
@@
Thus, on input $x$, $M^{O}_{0}$ has sufficient time to simulate $D^{O}_{x}$ and accepts if and only if $D^{O}_{x}$ rejects, this is also because in simulation of a polynomial-time deterministic oracle Turing machine with oracle $O$, $M^{O}_{0}$ only turns itself off mandatorily when the counter on tape $3$ overflows, i.e., when the counter
@@
We thus conclude from the above argument that $D^{O}_{i}$ does not exist in the enumeration $e$, i.e., $L^{O}_{d}$ not accepted by any machine in the enumeration $e$. In other words,
@@
By a similar argument appearing in the proof of Theorem 10, it is not hard to see that for any input $w$ to $M^{O}_{0}$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $M^{O}_{0}$ runs at most $|w|^{i_{w}+1}$ steps. Namely, $M^{O}_{0}$ runs within time
@@
Note further that for any fixed $i\in\mathbb{N}_{1}$, $L^{O}_{d,i}$ is the language accepted by a nondeterministic oracle Turing machine $M^{O}_{0}$ within $O(n^{i})$ steps, we thus obtain
(9) $\begin{split}L^{O}_{d,i}\in\text{NTIME}^{O}[n^{i}]\subseteq\mathcal{NP}^{O},\quad\text{for any fixed $i\in\mathbb{N}_{1}$}.\end{split}$
From (7), (8) and (9) we deduce that
@@
there is always a machine $D^{O}_{S}$ which is in $P^{O}$, such that
@@
for all $i\in\mathbb{N}_{1}$, thus contradicting the assumption that $P^{O}$ is enumerable.303030 The language accepted by machine $D^{O}_{S}$ differs from the languages accepted by all of the polynomial-time deterministic oracle Turing machines in the enumeration, but it lies in $P^{O}$ since $\mathcal{P}^{O}=\mathcal{NP}^{O}.$
The reason why for any enumeration of $P^{O}$ (when supposing $P^{O}$ is enumerable), there is always a machine $D^{O}_{S}$ in $P^{O}$ accepting the language
@@
The above arguments indicate that, we are unable to diagonalize against the set of $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$, just as Cantor [Can91] was unable to put all real numbers in the open interval $(0,1)$ into the slots indexed by all $i\in\mathbb{N}_{1}$.313131 For a more detailed comparison, the reader could consult the second proof (due to Cantor) that the continuum is not enumerable. See [Hob21], p. 82.
@@
does not qualify the technique of diagonalization by a universal nondeterministic oracle Turing machine with oracle $O$ to separate $\mathcal{P}^{O}$ and $\mathcal{NP}^{O}$. In other words, diagonalization techniques (via a universal nondeterministic oracle Turing machine) would not apply to the relativized versions of the $\mathcal{P}$ versus $\mathcal{NP}$ problem because of that $P^{O}$ is not enumerable in this case, as can be seen from above that we can always construct a machine $D^{O}_{S}$ such that, for any function
@@
Or equivalently, the cardinality of $P^{O}$ is greater than that of $\mathbb{N}_{1}$ (i.e., there is no $(1,1)$ correspondence between the set $P^{O}$ and $\mathbb{N}_{1}$), whereas the cardinality of $P$ is equal to that of $\mathbb{N}_{1}$ in Section 4, which is the most significant difference. In brief, the fact that the set $P$ of all polynomial-time deterministic Turing machines is enumerable is an important prerequisite for the application of diagonalization technique.
Moreover, even if suppose that $P^{O}$ is enumerable and $\mathcal{P}^{O}=\mathcal{NP}^{O}$, the diagonalization techniques (via a universal nondeterministic oracle Turing machine) would not apply as well, because $M_{0}^{O}$ is also in $P^{O}$ in this case and $M_{0}^{O}$ is unable to diagonalize against itself.
Remark 7.2
As the reader can see, we follow all of the Cantor’s mathematical premises of the diagonalizing approach when proving the Theorem 1 and Theorem 6. And, as a matter of fact, following all of the mathematical premises is exactly what we need to do when applying the diagonalizing technique.
8 Conclusions
To summarize, we have shown that there exists a language $L_{d}$ that is accepted by some nondeterministic Turing machines but by no polynomial-time deterministic Turing machines. To achieve this, we first encode any single-tape deterministic Turing machine into a positive integer (i.e., a binary string) by using the method given in [AHU74]. After that, the polynomial-time deterministic Turing machine $(M,k)$ could be the concatenation of the binary string representing $M$ itself followed by the order $110^{k}$ 1 of $(M,k)$. Our encoding of a polynomial-time deterministic Turing machine makes it very convenient for us to map any a polynomial-time deterministic Turing machine to a positive integer, thus showing that the set $P$ of all polynomial-time deterministic Turing machines is enumerable, and simultaneously, the encoding makes it very convenient for us to control the running time, i.e., the time complexity of the simulating machine $M_{0}$ constructed in Section 4.
Next, we design a four-tape universal nondeterministic Turing machine $M_{0}$ that diagonalizes against all of the polynomial-time deterministic Turing machines. The Theorem 9 illustrates the operation of the universal nondeterministic Turing machine $M_{0}$ in detail, showing that there is a language $L_{d}$ accepted by this universal nondeterministic Turing machine $M_{0}$ but by no polynomial-time deterministic Turing machines. In Theorem 10, we carefully analyse the running time of the universal nondeterministic Turing machine $M_{0}$, showing that it runs within time
@@ 8 Conclusions
Combining Theorem 9 and Theorem 21, Theorem 1 hence follows. As we observed, the techniques applied in this work, as an interesting application, can be utilized to show that one-way functions in the worst-case model (see p. 281–284 in [Pap94], or see also [Ko85, GS88]) do exist.
We have also presented a language $L_{s}$ which is in $\mathcal{P}$ and the machine accepting it also runs within time
@@
then the set $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$ is not enumerable. So the reader can convince himself that we cannot use the diagonalization technique by a universal nondeterministic oracle Turing machine with oracle $O$ to separate the complexity classes $\mathcal{P}^{O}$ and $\mathcal{NP}^{O}$ in the case of
@@
This shows that the so-called “Relativization Barrier” is not really a barrier but the polynomial-time oracle Turing machines, or the relativized versions of the $\mathcal{P}$ versus $\mathcal{NP}$ problem, are indeed with some mysterious.
@@
Are these two complexity classes the same? Also note that there is a subfield of computational complexity theory, namely, the proportional proof complexity, which was initiated by Cook and Reckhow [CR79] and is devoted to the goal of proving the conjecture
@@
We refer the reader to the reference [Coo00] for the importance of this research field and to the reference [Kra95] for the motivation of the development of this rich theory. Apart from this, the Chapter $10$ of [Pap94] also contains the introductions of importance of the problem
@@
Another intriguing open question is whether the cardinality of the set $P^{O}$ is less than that of real numbers if
$\mathcal{P}^{O}=\mathcal{NP}^{O}?\,$
Finally, although the today’s or future’s computers, which fall into the category of polynomial-time universal deterministic Turing machine (i.e., equivalent to the machine $M^{\prime}_{0}$ constructed in Section 6 for proving Theorem 12), cannot accurately calculate the entire $\mathcal{NP}$ set of problems, do not be pessimistic, our computers can solve as many problems as the cardinality of $\mathbb{N}_{1}$. Looking forward to discovering more and more practical polynomial-time deterministic algorithms to make our living environment much more convenient and much more colorful.
8.1 Acknowledgements
Sincere thanks go to complexity theorist Lance Fortnow [For21] for his valuable criticisms which cause us to seek Theorem 21.
Appendix A The Two Definitions of $\mathcal{P}$ (Respectively, $\mathcal{NP}$) Are Equivalent
The official descriptions on the definitions of $\mathcal{P}$ and $\mathcal{NP}$ in [Coo00] are given by:
Definition A.1
$\begin{array}[]{ll}\mathcal{P}\overset{{\rm def}}{=}&\{L\,|\,L=L(M)\text{ for some Turing machine $M$ that runs}\\ &\qquad\text{ in polynomial time }\}.\end{array}$
In Definition A.1, Turing machine $M$ refers to a deterministic Turing machine, and the notion of “running in polynomial time” is that if there exists $k$ ($\in\mathbb{N}_{1}$) such that,
$T_{M}(n)\leq n^{k}+k,\quad\forall n\in\mathbb{N}_{1}$
where $T_{M}(n)$ is defined by
$T_{M}(n)=\max\{t_{M}(w)|w\in\Sigma^{n}\}$
and $t_{M}(w)$ denotes the number of steps (or moves) in the computation of $M$ on input $w$ of length $n$.
Definition A.2
$\mathcal{NP}\overset{{\rm def}}{=}\{L\,|\,w\in L\Leftrightarrow\exists y(|y|\leq|w|^{k}\,\,\,\text{and }R(w,y))\}$
where the checking relation
$R(w,y)$
is polynomial-time, i.e., the language
$L_{R}\overset{{\rm def}}{=}\{w\#y|R(w,y)\}$
is in $\mathcal{P}$.
As we mentioned earlier, these two definitions look a little different from those given in Section 2. But they are equivalent. Here, we will show the equivalence for these two definitions of $\mathcal{P}$ and the equivalence for those two definitions of $\mathcal{NP}$.
A.1 A Proof of Two Definitions of $\mathcal{P}$ Are Equivalent.
Now, we are at an opportune point to show that the above two definitions of $\mathcal{P}$ are equivalent.
Proof. We show first the “if” part. Suppose that
$L\in\bigcup_{i\in\mathbb{N}_{1}}\text{DTIME}[n^{i}].$
Then there is a $k\in\mathbb{N}_{1}$, such that
$L\in\text{DTIME}[n^{k}]$
which means that for all $n\in\mathbb{N}_{1}$, there is a deterministic Turing machine $M$, for any $w\in\Sigma^{n}$
$T_{M}(|w|)\leq c_{0}n^{k}+c_{1}n^{k-1}+\cdots+c_{k-1}n+c_{k}\quad\text{where $c_{0}>0$}.$
and
$L=L(M).$
For any constants $c_{0}$, $c_{1}$, $\cdots$, $c_{k}$, there must exists a minimal $t\in\mathbb{N}_{1}$ such that for all $n\in\mathbb{N}$ and for any $w\in\Sigma^{n}$
$T_{M}(|w|)\leq n^{t}+t.$
So,
$L\in\{L\,|\,L=L(M)\text{ for some Turing machine $M$ that runs in polynomial time}\}.$
We show next the “only if” part. Suppose now that the language
$L\in\{L\,|\,L=L(M)\,\text{ for some Turing machine $M$ that runs in polynomialtime}\}.$
Then there exists a $k\in\mathbb{N}_{1}$ such that for all $n\in\mathbb{N}_{1}$ and for all $w\in\Sigma^{n}$,
$T_{M}(|w|)\leq n^{k}+k$
which implies that
$L\in\text{DTIME}[n^{k}]\subseteq\bigcup_{i\in\mathbb{N}_{1}}\text{DTIME}[n^{i}].$
The conclusion follows.
A.2 A Proof of Two Definitions of $\mathcal{NP}$ Are Equivalent.
The remainder of this Appendix is to show that Definition A.2 and the definition of $\mathcal{NP}$ given in Section 2 are equivalent.
To begin, we should first give another definition of $\mathcal{NP}$ which is similar to the official descriptions of the definition of $\mathcal{P}$ in [Coo00] (i.e., Definition A.1 above):
Definition A.3
$\begin{array}[]{ll}\mathcal{NP}\overset{{\rm def}}{=}&\{L\,|\,L=L(M)\text{ forsome nondeterministic Turing machine $M$}\\ &\qquad\text{that runs in polynomial time}\}.\end{array}$
Now, let us finish the last step of the task that we mentioned at the beginning of this subsection. Proof. We need to show that Definition A.3 is equivalent to Definition A.2 and to show the equivalence between Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2. By a similar argument given in Subsection A.1, it is clear that Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2 are equivalent, thus this part of proof will be omitted for brevity.
It is time for us to show that Definition A.3 is equivalent to Definition A.2. For the purposes of brevity, we omit the proof and refer the reader to [Kar72] (cf. Theorem 1 in [Kar72]), or to the proof of Theorem $7$.20202020 in [Sip13] which is a theorem answering the same question discussed here (see p. $294$ of [Sip13]). Thus this completes the proof.
Tianrong Lin
National Hakka University, China

### v33

@@ 7.1. Proof of Theorem 6
Since the input tape and the working tape of an oracle Turing machine can be the same tape, we thus assume that the machines in the set $P^{O}$ of the polynomial-time deterministic oracle Turing machines with oracle $O$ are two-tape oracle Turing machines; one is the input tape, and the other is the oracle tape.
Now, we are ready to prove Theorem 6:
@@
Next, we construct a five-tape universal nondeterministic oracle Turing machine $M_{0}^{O}$ that operates as follows on an input string $x$ of length $n$:
  1. (1)
    $M^{O}_{0}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some polynomial-time deterministic oracle Turing machine $D^{O}_{j}$ for some $j$, then GOTO $6$; else determine $t$, the number of tape symbols used by $D^{O}_{j}$; $s$, its number of states; and $k$, its order.292929 We suppose that the order of $D^{O}$, i.e., the minimal degree of some polynomial of $D^{O}$, is also encoded into $D^{O}$, similarly to the way presented in Section 3. The third tape of $D^{O}_{0}$ can be used as “scratch” memory to calculate $t$.
  2. (2)
@@
    cells each, the blocks being separated by a single cell holding a marker ##\##, i.e., there are
@@
  4. (4)
    On tape $4$, $M^{O}_{0}$ reads and writes the contents of the oracle tape of $D^{O}_{j}$. That is, tape $4$ is the oracle tape of $M^{O}_{0}$, which is used to simulate the oracle tape of $D^{O}_{j}$.
  5. (5)
@@
  6. (6)
    Since $x$ is not an encoding of some polynomial-time deterministic oracle Turing machine with oracle $O$. Then $M^{O}_{0}$ sets up a block of
@@
The nondeterministic oracle Turing machine $M^{O}_{0}$ described above is a nondeterministic oracle Turing machine that is of time complexity
@@
Suppose now $L^{O}_{d}$ were accepted by the $i$-th deterministic oracle Turing machine $D^{O}_{i}$ in the enumeration $e$, which is a deterministic
$T(n)=n^{k}+k$
time-bounded oracle Turing machine. Let $D^{O}_{i}$ have $s$ states and $t$ tape symbols. Since303030 See (III) of footnote $6$.
@@
So, there exists an $N_{0}>0$ such that for any $N\geq N_{0}$,
$T(N)\log T(N)<N^{k+1},$
@@
Thus, on input $x$, $M^{O}_{0}$ has sufficient time to simulate $D^{O}_{x}$ and accepts if and only if $D^{O}_{x}$ rejects. This is also because in the simulation of a polynomial-time deterministic oracle Turing machine with oracle $O$, $M^{O}_{0}$ only turns itself off mandatorily when the counter on tape $3$ overflows, i.e., when the counter
@@
We thus conclude from the above argument that $D^{O}_{i}$ does not exist in the enumeration $e$, i.e., $L^{O}_{d}$ is not accepted by any machine in the enumeration $e$. In other words,
@@
By a similar argument appearing in the proof of Theorem 4.4, it is not hard to see that for any input $w$ to $M^{O}_{0}$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $M^{O}_{0}$ runs at most $|w|^{i_{w}+1}$ steps. Namely, $M^{O}_{0}$ runs within time
@@
Similarly, the above definition of language $L^{O}_{d,i}$ technically still can be done by adding a new tape to $M_{0}^{O}$ as a counter to count up to
$n^{i+1}$
for a fixed $i\in\mathbb{N}_{1}$, meaning that $M_{0}^{O}$ turns itself off when the counter of tape $3$ exceeds
$n^{k+1}$
or the counter of the newly added tape exceeds
$n^{i+1}.$
@@
Note further that for any fixed $i\in\mathbb{N}_{1}$, $L^{O}_{d,i}$ is the language accepted by a nondeterministic oracle Turing machine $M^{O}_{0}$ within $O(n^{i})$ steps; we thus obtain
(9) $\begin{split}L^{O}_{d,i}\in{\rm NTIME}^{O}[n^{i}]\subseteq\mathcal{NP}^{O},\quad\text{for any fixed $i\in\mathbb{N}_{1}$}.\end{split}$
From (7), (8), and (9), we deduce that
@@
there is always a machine $D^{O}_{S}$ that is in $P^{O}$, such that
@@
for all $i\in\mathbb{N}_{1}$, thus contradicting the assumption that $P^{O}$ is enumerable.313131 The language accepted by machine $D^{O}_{S}$ differs from the languages accepted by all of the polynomial-time deterministic oracle Turing machines in the enumeration, but it lies in $\mathcal{P}^{O}$ since $\mathcal{P}^{O}=\mathcal{NP}^{O}.$
The reason why, for any enumeration of $P^{O}$ (when supposing $P^{O}$ is enumerable), there is always a machine $D^{O}_{S}$ in $P^{O}$ accepting the language
@@
The above arguments indicate that we are unable to diagonalize against the set of $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$, just as Cantor [Can91] was unable to put all real numbers in the open interval $(0,1)$ into the slots indexed by all $i\in\mathbb{N}_{1}$.323232 For a more detailed comparison, the reader could consult the second proof (due to Cantor) that the continuum is not enumerable. See [Hob21], p. 82.
@@
does not qualify the technique of diagonalization by a universal nondeterministic oracle Turing machine with oracle $O$ to separate $\mathcal{P}^{O}$ and $\mathcal{NP}^{O}$. In other words, diagonalization techniques (via a universal nondeterministic oracle Turing machine) would not apply to the relativized versions of the $\mathcal{P}$ versus $\mathcal{NP}$ problem because the set $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$ is not enumerable in this case, as can be seen from above that we can always construct a machine $D^{O}_{S}$ such that, for any function
@@
Or equivalently, the cardinality of $P^{O}$ is greater than that of $\mathbb{N}_{1}$ (i.e., there is no $(1,1)$ correspondence between the set $P^{O}$ and $\mathbb{N}_{1}$), whereas the cardinality of the set $P$ of all polynomial-time deterministic Turing machines is equal to that of $\mathbb{N}_{1}$ in Section 4, which is the most significant difference. In brief, the fact that the set $P$ of all polynomial-time deterministic Turing machines being enumerable is an important prerequisite for the application of the diagonalization techniques.
Moreover, even if we suppose that $P^{O}$ is enumerable and $\mathcal{P}^{O}=\mathcal{NP}^{O}$, the diagonalization techniques (via a universal nondeterministic oracle Turing machine) would not apply as well, because $M_{0}^{O}$ is also in $P^{O}$ in this case and $M_{0}^{O}$ is unable to diagonalize against itself.
Remark 7.3.
As the reader can see, we follow all of Cantor’s mathematical premises of the diagonalizing approach when proving Theorem 1.1 and Theorem 6. And, as a matter of fact, following all of the mathematical premises is exactly what we need to do when applying the diagonalization techniques.
8. Concluding Remarks and Open Problems
To summarize, we have shown that there exists a language $L_{d}$ that is accepted by some nondeterministic Turing machines but by no polynomial-time deterministic Turing machines. To achieve this, we first encode any single-tape deterministic Turing machine into a positive integer (i.e., a binary string) by using the method given in [AHU74]. After that, the polynomial-time deterministic Turing machine $(M,k)$ could be the concatenation of the binary string representing $M$ itself followed by the order $110^{k}$ 1 of $(M,k)$. Our encoding of a polynomial-time deterministic Turing machine makes it very convenient for us to map any polynomial-time deterministic Turing machine to a positive integer, thus showing that the set $P$ of all polynomial-time deterministic Turing machines is enumerable, and simultaneously, the encoding makes it very convenient for us to control the running time, i.e., the time complexity of the simulating machine $M_{0}$ constructed in Section 4.
Next, we design a four-tape universal nondeterministic Turing machine $M_{0}$ that diagonalizes against all of the polynomial-time deterministic Turing machines. The Theorem 4.1 illustrates the operation of the universal nondeterministic Turing machine $M_{0}$ in detail, showing that there is a language $L_{d}$ accepted by this universal nondeterministic Turing machine $M_{0}$ but by no polynomial-time deterministic Turing machines. In Theorem 4.4, we carefully analyze the running time of the universal nondeterministic Turing machine $M_{0}$, showing that it runs within time
@@ 8. Concluding Remarks and Open Problems
Combining Theorem 4.1 and Theorem 21, Theorem 1.1 hence follows. As we observed, the techniques applied in this work, as an interesting application, can be utilized to show that one-way functions in the worst-case model (see p. 281–284 in [Pap94], or see also [Ko85, GS88]) do exist.
We have also presented a language $L_{s}$ that is in $\mathcal{P}$, and the machine accepting it also runs within time
@@
then the set $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$ is not enumerable. So the reader can convince himself that we cannot use the diagonalization techniques by a universal nondeterministic oracle Turing machine with oracle $O$ to separate the complexity classes $\mathcal{P}^{O}$ and $\mathcal{NP}^{O}$ in the case of
@@
This shows that the so-called “relativization barrier” is not really a barrier, but the polynomial-time oracle Turing machines, or the relativized versions of the $\mathcal{P}$ versus $\mathcal{NP}$ problem, are indeed somewhat mysterious.
@@
Are these two complexity classes the same? Also note that there is a subfield of computational complexity theory, namely, proportional proof complexity, which was initiated by Cook and Reckhow [CR79] and is devoted to the goal of proving the conjecture
@@
We refer the reader to the reference [Coo00] for the importance of this research field and to the reference [Kra95] for the motivation of the development of this rich theory. Apart from this, Chapter $10$ of [Pap94] also contains the introductions of the importance of the problem
@@
Another intriguing open question is whether the cardinality of the set $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$ is less than that of real numbers if